\begin{thebibliography}{25}

\bibitem {banach} S. Banach and A. Tarski, Sur la decomposition des
ensembles de points en parties respectivement congruentes, \emph{{Fund.}
Math.}, 1924, vol.~6, pp.~244--277.

\bibitem{Kolm1} A.N. Kolmogorov, Three Approaches to the Quantitative
Definition of Information, \emph{{Probl.} Peredachi Inf.}, 1965, vol.~1,
no.~1, pp.~3--11 [\emph{{Probl.} {Inf.} Trans.} ({Engl.} Transl.),
pp.~1--7].

\bibitem{ku87} A.N. Kolmogorov and V.A. Uspensky, Algorithms and
Randomness, \emph{{Teor.} {Veroyatn.} Primen.}, 1987, vol.~32, no.~3,
pp.~425--455 [\emph{Theory {Probab.} Appl.} ({Engl.} Transl.),
pp.~389--412].

\bibitem {zl} Alexander K. Zvonkin and Leonid A. Levin, Complexity of
Finite Objects and the Algorithmic Concepts of Information and
Randomness, \emph{{Usp.} {Mat.} Nauk}, 1970, vol.~25, no.~6, pp.~85--127
[\emph{Russian {Math.} Surveys\/} ({Engl.} Transl.), pp.~83--124].

\bibitem {RSA} R.L. Rivest, A. Shamir, and L.M. Adleman, A Method for
Obtaining Digital Signatures and Public-Key Cryptosystems,
\emph{{Commun.} ACM}, 1978, vol.~21, no.~2, pp.~120--126.

\bibitem{K67} A.N. Kolmogorov, Several Theorems about Algorithmic
Entropy and Algorithmic Amount of Information (a talk at a Moscow
{Math.} {Soc.} meeting 10/31/67). An abstract in \emph{{Usp.} {Mat.}
Nauk}, 1968, vol.~23, no.~2, p.~201.

\bibitem {BM} M. Blum and S. Micali, How to Generate Cryptographically
Strong Sequences of Pseudo-Random Bits, \emph{SIAM J. Comp.}, 1984,
vol.~13, pp.~850--864.

\bibitem {yao} A.C. Yao, Theory and Applications of Trapdoor Functions,
in \emph{{Proc.} of the 23rd {Ann.} IEEE {Sympos.} on Foundations of
Computer Science}, 1982, pp.~80--91.

\bibitem{ow} Leonid A. Levin, One-Way Functions and Pseudorandom
Generators, \emph{Combinatorica}, 1987, vol.~7, no.~4, pp.~357--363.

\bibitem{GL} Oded Goldreich and Leonid A. Levin, A Hard-Core Predicate
for any One-way Function, in \emph{{Proc.} of the 21st {Ann.} ACM
{Sympos.} on Theory of Computing}, 1989, pp.~25--32.

\bibitem{L93} Leonid A. Levin, Randomness and Non-determinism, \emph {J.
{Symb.} Logic}, 1993, vol.~58, no.~3, pp.~1102--1103.\\
 A less technical version is in \emph {International Congress of
Mathematicians}, Zurich, August 1994. \emph {Proceedings} (Invited
Addresses), pp. 1418--1419. Birkhauser Verlag, 1995.

\bibitem{Knuth} Donald E. Knuth, \emph{The Art of Computer Programming},
vol.~2: \emph{Seminumerical Algorithms}, Reading, Mass.: Addison-Wesley,
1997, 3rd ed., Sec.~3.5.F.

\bibitem{HILL} Johan Hastad, Russell Impagliazzo, Leonid A. Levin, and
Michael Luby, A Pseudorandom Generator from any One-way Function,
\emph{SIAM J. Comp.}, 1999, vol.~28, no.~4, pp.~1364--1396.

\bibitem{NY} M. Naor and M. Yung, Universal One-way Hash Functions and
Their Applications, in \emph{{Proc.} of the 21st {Ann.} ACM {Sympos.} on
Theory of Computing}, 1989, pp.~33--43.

\bibitem{GMW} O. Goldreich, S. Micali, and A. Wigderson, Proofs That
Yield Nothing but Their Validity, \emph{J. ACM}, 1991, vol.~38, no.~3,
pp.~691--729.

\bibitem {shamir} A. Shamir, Factoring Numbers in $O(\log n)$ Arithmetic
Steps, \emph{{Inf.} {Process.} Lett.}, 1979, vol.~8, no.~1, pp.~28--31.

\bibitem {shor} P.W. Shor, Polynomial-Time Algorithms for Prime
Factorization and Discrete Logarithms on a Quantum Computer, \emph{SIAM
J. Comp.}, 1997, vol.~26, no.~5, pp.~1484--1509.

\bibitem {shor-spr} P.W. Shor, ``Re: When will quantum computers become
practical?''\\ Usenet group \emph{sci.physics.research}, 2000/03/22.
Available from\\
\verb" http://www.google.com/groups?threadm=FrsHL0.ILr%40research.att.com "

\bibitem{rp} Leonid A. Levin, Average Case Complete Problems, \emph{SIAM
J. Comp.}, 1986, vol.~15, no.~1, pp.~285--286.

\bibitem{samp} Russell Impagliazzo and Leonid A. Levin, No Better Ways
to Generate Hard NP Instances than Picking Uniformly at Random, in
\emph{{Proc.} of the 31st {Ann.} IEEE {Sympos.} on Foundations of
Computer Science}, 1990, pp.~812--821.

\bibitem{search} Leonid A. Levin, Universal Search Problems,
\emph{{Probl.} Peredachi Inf.}, 1973, vol.~9, no.~3, pp.~115--116
[\emph{{Probl.} {Inf.} Trans.} ({Engl.} Transl.), pp.~265--266].

\bibitem{laws} Leonid A. Levin, Randomness Conservation Inequalities,
\emph{{Inf.} Control}, 1984, vol.~61, no.~1, pp.~15--37.

\bibitem{bnt} C.H. Bennett, Logical Depth and Physical Complexity,
\emph{The Universal Turing Machine---A Half-Century Survey}, Herken, R.,
Ed., Oxford: Oxford {Univ.} Press, 1988, pp.~227--257.

\bibitem{color} Ramarathnam Venkatesan and Leonid A. Levin, Random
Instances of a Graph Coloring Problem Are Hard, in \emph{{Proc.} of the
20th {Ann.} ACM {Sympos.} on Theory of Computing}, 1988, pp.~217--222.

\bibitem{G90} Yu. Gurevich, Matrix Decomposition Is Complete for the
Average Case, in \emph{{Proc.} of the 31st {Ann.} IEEE {Sympos.} on
Foundations of Computer Science}, 1990, pp.~802--811.

\bibitem{VR} R. Venkatesan and S. Rajagopalan, Average Case
Intractability of Matrix and Diophantine Problems, in \emph{{Proc.} of
the 24th {Ann.} ACM {Sympos.} on Theory of Computing}, 1992,
pp.~632--642.

\bibitem{JW} J. Wang, Average Case Completeness of a Word Problem for
Groups, in \emph{{Proc.} of the 27th {Ann.} ACM {Sympos.} on Theory of
Computing}, 1995, pp.~325--334.

\bibitem{expnd} Oded Goldreich, Russell Impagliazzo, Leonid A. Levin,
Ramarathnam Venkatesan, and David Zuckerman, Security Preserving
Amplification of Hardness, in \emph{{Proc.} of the 31st {Ann.} IEEE
{Sympos.} on Foundations of Computer Science}, 1990, pp.~318--326.

\end{thebibliography}

\documentclass[12pt,twoside]{article}\textheight 620pt\textwidth 6.5in
 \usepackage{amssymb} 
\begin{document}

 \newtheorem{conjecture}{Conjecture} \newtheorem{lemma}{Lemma}
 \newtheorem{notation}{Notation} \newtheorem{theorem}{Theorem}
 \newtheorem{definition}{Definition} \newtheorem{remark}{Remark}
 \newtheorem{proposition}{Proposition}

 \title{The Tale of One-Way Functions\thanks
 {Problemy Peredachi Informatsii 39(1), 2003 \
 ( = Problems of Information Transmission),\hfil\mbox{}
 UDK~621.391:519.2}}

 \author{Leonid A. Levin\thanks
        {This research was partially conducted by the author
 for the Clay Mathematics Institute and supported by the Institut
 des Hautes \'Etudes Scientifiques and NSF, grant no.~CCR-9820934.}\\
 Boston University\thanks
 {Computer {Sci.} dept., 111 Cummington St., Boston, MA 02215.}}

 \date{} \maketitle \begin{flushright}\begin{tabular}{l}
 \emph {All the king's horses, and all the king's men,}\\*
 \emph {Couldn't put Humpty together again.}
 \end{tabular}\end{flushright}

\begin{abstract} The existence of one-way functions (owf) is arguably
the most important problem in computer theory. The article discusses and
refines a number of concepts relevant to this problem. For instance, it
gives the first combinatorial complete owf, i.e., a function which is
one-way if any function is. There are surprisingly many subtleties in
basic definitions. Some of these subtleties are discussed or hinted at
in the literature and some are overlooked. Here, a~unified approach is
attempted. \end{abstract}

\section{Introduction I: Inverting Functions}

From time immemorial, humanity has gotten frequent, often cruel,
reminders that many things are easier to do than to reverse. When the
foundations of mathematics started to be seriously analyzed, this
experience immediately found a formal expression.

\subsection{An Odd Axiom}

Over a century ago, George Cantor reduced all the great variety of
mathematical concepts to just one---the concept of sets---and derived
all mathematical theorems from just one axiom scheme---\emph{Cantor's
Postulate}. For each set-theoretical formula $A(x)$, it postulates the
existence of a set containing those and only those $x$ satisfying $A$.
This axiom looked a triviality, almost a definition, but was soon found
to yield more than Cantor wanted, including contradictions. To salvage
its great promise, Zermelo, Fraenkel, and others pragmatically replaced
Cantor's Postulate with a collection of its restricted cases, limiting
the types of allowed properties $A$. The restrictions turned out to
cause little inconvenience and precluded (so far) any contradictions;
the axioms took their firm place in the foundations of mathematics.

In 1904, Zermelo noticed that one more axiom was needed to derive all
known mathematics, the (in)famous Axiom of Choice: every function $f$
has an inverse $g$ such that $f(g(x))=x$ for $x$ in the range of $f$. It
was accepted reluctantly; to this day, proofs dependent on it are being
singled out. Its strangeness was not limited to going beyond Cantor's
Postulate---it brought paradoxes! Allow me a simple illustration based
on the ability of the axiom of choice to enable a \emph{symmetric}
choice of an arbitrary integer.

Consider the additive group $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ of reals
mod $1$ as points $x\in[0,1)$ on a circle; take its subgroup
$Q_{10}\subset\mathbb{T}$ of decimal fractions $a/10^b$. Let $f(x)$ be
the (countable) coset $x+Q_{10}$, i.e., $f$ projects $\mathbb{T}$ onto
its factor group $\mathbb{T}/Q_{10}$. Any inverse $g$ of $f$ then
selects one representative from each coset. Denote by
$G=g(f(\mathbb{T}))$ the image of such a $g$; then each $x\in\mathbb{T}$
is brought into $G$ by exactly one rational shift $x+q$, $q\in Q_{10}$.
Now I will deviate from the standard path to emphasize the elementary
nature of the paradox. One last notation: $q'= (10 q)\bmod 1$ is $q\in
Q_{10}$, shortened by the removal of its most significant digit.

For a pair $p,q\in Q_{10}$, I bet $2:1$ that $x+p$ rather than $x+q$
falls in $G$ for a random $x\in\mathbb{T}$. The deal should be
attractive to you since my bet is higher while conditions to win are
completely symmetric under rotation of $x$. To compound my charitable
nature to its extreme, I offer such bets for all $q\in Q_{10}$, $q>0$,
$p=q'$, not just one pair. If you rush to accept, we choose a random $x$
(by~rolling dice for all of its digits) and find the unique ${\bf q}\in
Q_{10}$ for which $x+{\bf q}\in G$. Then I lose one bet for this ${\bf
q}$ and win ten (for each $q$ such that $q'={\bf q}$). Generosity pays!

This paradox is not as easy to dismiss as is often thought. Only 11 bets
are paid in each game: no infinite pyramids. Moreover, if $x$ is drawn
from a sphere $S_2$, a finite number of even unpaid bets suffices:
Banach and Tarski \cite{banach} constructed 6 pairs, each including a
set $A_i\subset S_2$ and a rotation~$T_i$; betting $x\in A_i$ versus
$T_i(x)\in A_i$, they lose one bet and win two for each $x$. Our $x+p$
and $x+q$ above are tested for the same condition and differ in finitely
many digits; all digits are evenly distributed and independent. One can
refuse the thought experiment of rolling the infinite number of digits
of $x$ or the question of whether $x+q\in G$, but this amounts to
rejecting basic concepts of the set theory. It is simpler to interpret
the refusal to bet as a hidden disbelief in the Axiom of Choice.

\subsection{Finite Objects: Exhaustive Search}

These problems with inverting functions have limited relevance to
computations. The latter deal with finite objects, which are naturally
well-ordered by Induction Axioms, rendering the Axiom of Choice
unnecessary. There are other mitigating considerations, too. Shannon
Information theory assigns to a random variable $x$ as much information
about its function $f(x)$ as to $f(x)$ about $x$. Kolmogorov
(algorithmic) information theory (see~\cite{Kolm1,ku87}) extended this
concept from random to arbitrary $x$: $I(x:y)$ is the difference between
the smallest lengths of programs generating $y$ and of those
transforming $x$ to $y$. Kolmogorov and I proved in 1967 that this
measure is symmetric too, like Shannon's, albeit approximately
\cite{zl}.

The proof involved a caveat: computationally prohibitive exhaustive
search of all strings of a given length. For instance, the product $pq$
of two primes contains as much (i.e., all) information about them as
vice versa, but \cite{RSA} and great many other things in modern
cryptography depend on the assumption that recovering the factors of
$pq$ is infeasible. Kolmogorov suggested at the time~\cite{K67} that
this information symmetry theorem may be a good test case to prove that
for some tasks exhaustive search cannot be avoided (in today's terms,
${\rm P}\ne{\rm NP}$).

The RSA application marked a dramatic twist in the role of the inversion
problem: from notorious troublemaker to priceless tool. RSA was the
first of the myriad bewildering applications which soon followed. At the
heart of many of them was the discovery that the hardness of every
one-way function $f(x)$ can be focused into a \emph{hard-core} bit,
i.e., an easily computable predicate $b(x)$ which is as hard to
determine from $f(x)$, or even to guess with any noticeable correlation,
as to recover $x$ completely.

In \cite{BM}, the first hard-core for $f(x)= a^x\bmod p$ was found,
which was soon followed with hard-cores for RSA and Rabin's function
$(x^2\bmod n)$, $n=pq$, and for ``grinding'' functions
$f^*(x_1,\ldots,x_n)= f(x_1),\ldots, f(x_n)$ (see \cite{yao}; for the
proof of Isolation Lemma used in \cite{yao}, see \cite{ow}). In
\cite{GL}, the general case was proved (see also \cite {L93,Knuth}).

The importance of such hard-cores comes from their use, proposed in
\cite{BM,yao} for unlimited deterministic generation of perfectly random
bits from a small random seed $s$. In the case of permutation $f$, such
generators are straightforward: $g_s(i)=b(f^i(s))$,
$i=0,1,2,\ldots\strut$. The general case was worked out in \cite{HILL}.

With the barrier between random and deterministic processes thus broken,
many previously unthinkable feats were demonstrated in the 80s. Generic
cryptographic results (such as, e.g., \cite{NY}), zero-knowledge proofs,
and implementation of arbitrary protocols between distrusting parties
(e.g., card games) as full information games \cite{GMW} are just some of
the many famous examples. This period was truly a golden age of Computer
Theory brought about by the discovery of the use of one-way functions.

\section{Introduction II: Extravagant Models}

\subsection{The Downfall of RSA}

This development was all the more remarkable as the very existence of
one-way (i.e., easy to compute, infeasible to invert) functions remains
unproven and subject to repeated assaults. The first came from Shamir
himself, one of the inventors of the RSA system. He proved \cite{shamir}
that factoring (on infeasibility of which RSA depends) can be done in
polynomial number of arithmetic operations. This result uses a so-called
``unit-cost model,'' which charges one unit for each arithmetic
operation, however long the operands. Squaring a number doubles its
length, repeated squaring brings it quickly to cosmological sizes.
Embedding a huge array of ordinary numbers into such a long one allows
one arithmetic step to do much work, e.g., to check exponentially many
factor candidates. The closed-minded cryptographers, however, were not
convinced and this result brought a dismissal of the unit-cost model,
not RSA.

Another, not dissimilar, attack is raging this very moment. It started
with the brilliant result of Peter Shor. He factors integers in
polynomial time using an imaginary analog device, Quantum Computer (QC),
inspired by the laws of quantum physics taken to their formal extreme.

\subsection{Quantum Computers}

QC has $n$ interacting elements, called q-bits. A pure state of each is
a unit vector on the complex plane $\mathbb{C}^2$. Its two components
are quantum amplitudes of its two Boolean values. A state of the entire
machine is a vector in the tensor product of $n$ planes. Its $2^n$
coordinate vectors are tensor-products of q-bit basis states, one for
each $n$-bit combination. The machine is cooled, isolated from the
environment nearly perfectly, and initialized in one of its basis states
representing the input and empty memory bits. The computation is
arranged as a sequence of perfectly reversible interactions of the
q-bits, putting their combination in the superposition of a rapidly
increasing number of basis states, each having an exponentially small
amplitude. The environment may intervene with errors; the computation is
done in an error-correcting way, immune to such errors as long as they
are few and of special restricted forms. Otherwise, the equations of
Quantum Mechanics are obeyed with unlimited precision. This is crucial
since the amplitudes are exponentially small and deviations in remote
(hundredth or even much further) decimal places would overwhelm the
content completely. In \cite{shor}, such computers are shown capable of
factoring in polynomial time. The exponentially many coordinates of
their states can, using a rough analogy, explore one potential solution
each and concentrate the amplitudes in the one that works.

\subsection{Small Difficulties}

There are many problems with such QCs. For instance, thermal isolation
cannot be perfect. Tiny backgrounds of neutrinos, gravitational waves,
and other exotics, cannot be shielded. Their effects on quantum
amplitudes need not satisfy the restrictions on which error-correcting
tools depend. Moreover, nondissipating computing gates, even classical,
remain a speculation. Decades past, their existence was cheerfully
proclaimed and even proved for worlds where the laws of physical
interaction can be custom-designed. In our world, where the
electromagnetic interaction between electrons, nuclei, and photons is
about the only one readily available, circuits producing less entropy
than computing remain hypothetical. So, low temperatures have limits and
even a tiny amount of heat can cause severe decoherence problems.
Furthermore, the uncontrollable degrees of freedom need not behave
simply as heat. Interaction with the intricately correlated q-bits may
put them in devilish states capable of conspiracies which defy
imagination.

\subsection{Remote Decimals}

All such problems, however, are peanuts. The major problem is the
requirement that basic quantum equations hold to multi-hundredth if not
millionth decimal positions where the significant digits of the relevant
quantum amplitudes reside. We have never seen a physical law valid to
over a dozen decimals. Typically, every few new decimal places require
major rethinking of most basic concepts. Are quantum amplitudes still
complex numbers to \emph{such} accuracies or do they become quaternions,
colored graphs, or sick-humored gremlins? I suspect physicists would
doubt even the laws of arithmetic pushed that far. In fact, we know that
the most basic laws cannot all be correct to hundreds of decimals: this
is where they stop being consistent with each other!

And what is the physical meaning of 500 digit long numbers? What could
one possibly mean by saying ``This box has a remarkable property: its
many q-bits contain the Ten Commandments with the amplitude whose first
500 decimal places end with 666''? What physical interpretation could
this statement have even for just this one amplitude? Close to the
tensor product basis, one might have opportunities to restate the
assertions using several short measurable numbers instead of one long.
Such opportunities may also exist for large systems, such as lasers or
condensates, where individual states matter little. But QC factoring
uses amplitudes of an exponential number of highly individualized basis
states. I doubt anything short of the most generic and direct use of
these huge precisions would be easy to substitute. One can make the
amplitudes more ``physical'' by choosing a less physical basis. Let us
look into this.

\subsection{Too Small Universe}\label{small-u}

QC proponents often say they win either way, by making a working QC or
by finding a correction to Quantum Mechanics; e.g., in \cite{shor-spr}
Peter Shor said: ``If there are nonlinearities in quantum mechanics
which are detectable by watching quantum computers fail, physicists will
be VERY interested (I would expect a Nobel prize for conclusive evidence
of this).''

Consider, however, this scenario. With few q-bits, QC is eventually made
to work. The progress stops, though, long before QC factoring starts
competing with pencils. The QC people then demand some noble prize for
the correction to the Quantum Mechanics. But the committee wants more
specifics than simply a nonworking machine, so something like observing
the state of the QC is needed. Then they find the Universe too small for
observing individual states of the needed dimensions and accuracy.
(Raising sufficient funds to compete with pencil factoring may justify a
Nobel Prize in Economics.)

Let us make some calculations. In cryptography, the length $n$ of the
integers to factor may be a thousand bits (and could easily be
millions.) By $\sim n$, I will mean a reasonable power of $n$.
A~$2^{\sim n}$~dimensional space $H$ has $2^{2^{\sim n}}$ nearly
orthogonal vectors. Take a generic $v\in H$. The minimal size of a
machine which can recognize or generate $v$ (approximately) is
$K=2^{\sim n}$---far larger than our Universe. This comes from a
cardinality argument: $2^{\sim K}$ machines of $K$ atoms. Let us call
such $v$ ``megastates.''

There is a big difference between untested and untestable regimes.
Claims about individual megastates are untestable. I can imagine a
feasible way to separate any \emph{two} QC states \emph{from each
other}. However, as this calculation shows, no machine can separate a
generic QC state from the set of all states more distant from it than QC
tolerates. So, what thought experiments can probe the QC to be in the
state described with the accuracy needed? I would allow to use the
resources of the entire Universe, but \emph{not more}!

Archimedes made a great discovery that digital representation of numbers
is exponentially more efficient than analog ones (sand pile sizes). Many
subsequent analog devices yielded unimpressive results. It is not clear
why QCs should be an exception.

\subsection{Metric versus Topology}\label{metr}

A gap in quantum formalism may be contributing to the confusion.
Approximation has two subtly different aspects: metric and topology.
Metric tells how close our ideal point is to a specific wrong one.
Topology tells how close it is to the combination of all unacceptable
(nonneighboring) points. This may differ from the distance to the
closest unacceptable point, especially for quantum systems.

In infinite dimensions, the distinction between 0 and positive
separation between a point and a set varies with topologies. In finite
dimensions, 0-vs.-positive distinction is too coarse: all topologies
agree. Since $2^{500}$ is finite only in a very philosophical sense, one
needs a quantitative refinement, some sort of a weak-topological (not
metric) \emph{depth} of a neighborhood polynomially related to resources
required for precision to a given depth. Then, precision to reasonable
depths would be physical, e.g., allow one to generate points inside the
neighborhood, distinguish its center from the outside, etc.

Metric defines $\varepsilon$-neighborhoods and is richer in that than
topology where the specific value of $\varepsilon$ is lost (only
$\varepsilon>0$ is assured). However, metric is restricted by the axiom
that the intersection of any set of $\varepsilon$-neighborhoods is
always another $\varepsilon$-neighborhood. Quantum proximity may require
both benefits: defined depth $\varepsilon$ and freedom to express it by
formulas violating the ``intersection axiom.'' Here is an example of
such a violation, without pretense of relevance to our needs. Suppose a
neighborhood of 0 is given by a set of linear inequalities $f_i(x)<1$;
then its depth may be taken as $1/\sum\limits_i\|f_i\|$. Restricting $x$
to the unit sphere would render this depth quadratically close to metric
depth. A more relevant formula may need preferred treatment of tensor
product basis.

\subsection{The Cheaper Boon}

QC of the sort that factors long numbers seems firmly rooted in science
fiction. It is a pity that popular accounts do not distinguish it from
much more believable ideas, like Quantum Cryptography, Quantum
Communications, and the sort of Quantum Computing that deals primarily
with locality restrictions, such as fast search of long arrays. It is
worth noting that the reasons why QC must fail are by no means clear;
they merit thorough investigation. The answer may bring much greater
benefits to the understanding of basic physical concepts than any
factoring device could ever promise. The present attitude is analogous
to, say, Maxwell selling the Daemon of his famous thought experiment as
a path to cheaper electricity from heat. If he did, much of insights of
today's thermodynamics might be lost or delayed.

The rest of the article ignores any extravagant models and stands fully
committed to the Polynomial Overhead Church--Turing Thesis: Any
computation that takes $t$ steps on an $s$-bit device can be simulated
by a Turing Machine in $s^{O(1)}t$ steps within $s^{O(1)}$ cells.

\section{The Treacherous Averaging}

Worst-case hardness of inverting functions may bring no significant
implications. Imagine that all instances come in two types: ``easy'' and
``hard.'' The easy instances $x$ take $\|x\|^2$ time. An~exponential
expected time of randomized algorithms is required \emph{both} to solve,
and \emph{to find} any hard instance. So, the Universe would be too
small to ever produce intractable instances, and the inversion problem
would pose no practical difficulty. It is ``generic,'' not worst-case,
instances that both frustrate algorithm designers and empower
cryptographers to do their incredible feats. The definition of
``generic,'' however, requires great care.

\subsection{Las Vegas Algorithms}\label{las}

First we must agree on how to measure the performance of inverters.
Besides instances $x=f(w)$, algorithms $A(x,\alpha)$ inverting one-way
functions $f$ can use random dice sequences
$\alpha\in\{0,1\}^{\mathbb{N}}$. They never need a chance for (always
filterable) wrong answers. So we restrict ourselves to \emph{Las Vegas}
algorithms which can only produce a correct output, abort, or diverge.

For a given $x$, the performance of $A$ has two aspects: the
volume\footnote {I say \emph{volume} rather than \emph{time}, for
greater robustness in the case of massively parallel models.}
 $V$ of computation and the chance $p_V$ (over $\alpha$) of success in
$V$ steps. They are not independent: $p$ can be always boosted while
roughly preserving $V/p$ (more precisely $-V/\log(1\!-\!p)$)
 by simply running $A$ on several independent $\alpha$. This idea
suggests the popular requirement that Las Vegas algorithms be normalized
to, say, $p\ge \frac12$. The problem with this restriction is that
estimating $p$ and the needed number of trials may require exponential
volume overhead in the worst case. Thus, only such measures as average
volume can be kept reasonable while normalizing the chance. It is
important that both are averaged only over $A$'s own dice $\alpha$; the
instance $x$ is chosen by the adversary. In this setting, the duplicity
of performance aspects does become redundant: $p$ and average volume are
freely interchangeable (to shrink the latter, one simply runs $A$ with a
small chance).

Combining several runs into one lessens the modularity and counting the
runs needed does involve some overhead. However, there are more
substantial reasons to prefer normalization of average volume to that of
the success rate $p$. In some settings, success is a matter of degree.
For instance, different inverses of the same instance of a owf may be of
different and hard-to-compare value. Normalizing the average volume, on
the other hand, is robust. This volume bound may be $O(1)$ if the model
of computation is very specific. If flexibility between several
reasonable models is desired, polynomial bounds, specific to each
algorithm, may be preferable. There is one obstacle: the set of
algorithms with a restricted expected complexity is not recursively
enumerable. We can circumvent this problem by using the following
enforceable form of the bound.

\begin{definition}
\emph{Las Vegas} algorithms $A(x,\alpha)\in\mathop{\rm LV}(b)$ start
with a given bound $b(x)$ on expected computation volume. At any time,
the algorithm can bet a part of the remaining volume, so that it is
doubled or subtracted depending on the next dice of $\alpha$.
$\mathop{\rm LV}(O(1))$ usually suffices and we will abbreviate it as
L.\footnote{Pronounced ``Las'' algorithms to hint at Las Vegas and the
term's inventor Laszlo Babai. I would like to stress that no YACC (Yet
Another Complexity Class) is being introduced here. L is a \emph{form}
of algorithms; this is much less abstract than a class of algorithms or,
especially, a class of problems solvable by a class of algorithms.
Besides, it is not really new, just a slight tightening of the Las Vegas
restriction.} \end{definition}

Despite its tight $O(1)$ expected complexity bound, L is robust since
any Las Vegas algorithm can be put in this form, roughly preserving the
ratio between the complexity bound and the success rate. The inverse of
the latter gives the number of runs required for a constant chance of
success, thus playing the role of running time. An extra benefit is that
a reader adverse to bothering with the inner workings of computers can
just accept their restriction to L and view all further analysis in
purely probabilistic terms!

\subsection{Multimedian Time}\label{mult}

Averaging over the instance $x$ is, however, much trickier. It is not
robust to define generic complexity of an algorithm $A(x)$ running in
$t(x)$ steps as its expected time ${\bf E}_x t(x)$. A different device
may have a quadratic time overhead. For instance, recognition of input
palindromes requires quadratic time on a Turing machine with one tape,
but only linear time with two tapes. It may be that a similar overhead
exists for much slower algorithms, too. Then $t(x)$ may be, say,
$\|x\|^2$ for $x\not\in0^*$, while $t(0\ldots0)$ may be $2^{\|x\|}$ for
one device and $4^{\|x\|}$ for another. Take $x$ uniformly distributed
on $\{0,1\}^n$. Then ${\bf E}_x t(x)$ for these devices would be
quadratic and exponential respectively: averaging does not commute with
squaring. Besides, this exponential average hardness is misleading since
the hard instances would never appear in practice!

More device-independent would be the \emph{median} time, the minimal
number of steps spent for the \emph{harder half} of instances. This
measure, however, is not robust in another respect: it can change
dramatically if its \emph{half} threshold is replaced with, say, a
\emph{quarter}.

Fortunately, these problems disappear as one of the many benefits of our
Las Vegas conventions. One can simply take algorithms in L and measure
their chance of success for inputs chosen randomly with a given
distribution. The inverse of this chance, as a function of, say, input
length, is a robust measure of \emph{security} of a one-way function.
This measure is important in cryptography, where any noticeable chance
of breaking the code must be excluded. A different measure is required
for positive tasks aimed at success for almost all instances. We start
by considering a combined distribution over all instance lengths.

\begin{definition}
We consider an L-distribution of instances, i.e., a distribution of
outputs of an L-algorithm's on empty input.\footnote{If the instance
generator is not algorithmic, the definition can be modified to use the
output length instead of complexity in the definition of L. The
instances of length $n$ should have probabilities combining to a
polynomial, e.g., $p(x)=1/(\|x\|\log\|x\|)^2$.} Now, we run the
generator $k$ times (spending an average time of $O(k)$) and apply the
inverter until all generated solvable instances are solved.\footnote{If
the generator can produce unsolvable instances too, the definition
ignores them.} The number of trials is a random variable depending on
the inverter's and generator's dice. Its median value MT$(k)$ is called
the \emph{multimedian} time of inverting $f$ by $A$. \end{definition}

This measure is robust in many respects. It commutes with squaring the
inverter's complexity and, thus, is robust against variation of models.
It does not depend much on the $\frac12$ probability cut-off used for
median. Indeed, increasing $k$ by a factor of $c$ raises MT$(k)$ as much
as does tightening the inverter's failure probability to $2^{-c}$.

MT is relevant for both upper and lower bounds. Let T$(x)$ be high for
$\varepsilon$ fraction of $x\in\{0,1\}^n$. Then MT$(k)$ is as high for
$k=n^3/\varepsilon$. Conversely, let MT$(k)$ be high. Then, with
overwhelming probability, T$(x_i)$ is at least as high for some of
$n=k^2$ random $x_1,\ldots, x_n$ (and $\sum\limits_i\|x_i\|=O(n)$).

\subsection{Nice Distributions}\label{round}

So far, we addressed the variance of performance of a randomized
algorithm over its variable dice for a fixed input, as well as the issue
of averaging it over variable input with a given distribution. Now we
must address the variance of distributions. Choosing the right
distributions is not always trivial, a fact often dismissed by declaring
them uniform. Such declarations are confusing, though, since many
different distributions deserve the name.

For instance, consider graphs $G=(V,E\!\subset\! V^2)$, $\|V\|=n$, where
$n$ is chosen with probability $c/n^2$ (here, $c$ is a normalizing
factor). For a given $n$, graphs $G$ are chosen with two distributions,
both with a claim to uniformity: $\mu_1$ chooses $G$ with equal
probability among all $2^{n^2}$ graphs; $\mu_2$ first chooses $k=\|E\|$
with uniform probability $1/(n^2+1)$ and then $G$ with equal probability
among all the $C^k_{n^2}$ candidates. The set $\{G:\: \|E\|=n^{1.5}\}$
has then $\mu_2$ probability $1/(n^2+1)$, while its $\mu_1$ is
exponentially small. In fact, all nice distributions can be described as
uniform in a reasonable representation. Let me reproduce the argument,
sketched briefly in \cite{rp}, adding an additional aspect that I will
use later.

Let us use set-theoretic representation of integers: $n=\{0,1,\ldots,
n-1\}$. A measure $\mu$ is an additive real function of \emph{sets} of
integers; $\mu(n)=\mu(\{0\})+\mu(\{1\})+\ldots+\mu(\{n-1\})$ is its
monotone \emph{distribution} function. Its \emph{density} $\mu'(n)=
\mu(n+1)-\mu(n)=\mu(\{n\})$ is the probability of $\{n\}$ as a
singleton, rather than of a set $n=\{0,1,\ldots,n-\!1\}$. Let $Q_2$ be
the set of binary fractions $i/2^{\|i\|}\in[1/2,1)$. We round the
real-valued $\mu$ to $Q_2$, keeping only as many binary digits as needed
for constant factor accuracy of probabilities.

\begin{definition}
$\mu\colon\mathbb{N}\to Q_2$ is \emph{perfectly rounded} if $\mu(x)$ is
the shortest fraction within the interval
$\bigl(\mu(x-1),\mu(x+1)\bigr)$ and $-\log\mu(\{x\})=O(\|x\|)$.
\end{definition}

The last condition is just a convenience and can be met simply by mixing
the (monotone) $\mu$ with some simple distribution.

\begin{lemma}\label{rnd}
Every computable $\mu\colon\mathbb{N}\to Q_2$ can be uniformly
transformed into a perfectly rounded $\mu_1$ computable at most $\|x\|$
times slower than $\mu(x)$ and so that, for nondecreasing $\mu$ (i.e.,
one with $\mu'\ge 0$), $\mu'_1\ge\mu'/4$. \end{lemma}

Monotonicity is assured by comparing $\mu(x)$ with $\mu(y)$, for
prefixes $y$ of $x\in Q_2$. Then the claim can be achieved by rounding.
First, round $\mu(x)$ to the shortest binary $p$ that is closer to it
than to any other $\mu(y)$ and call these rounded values \emph{points}.
Find all \emph{slots}, i.e., closest to $p$ shorter binary fractions of
each binary length. Then, for each slot in the order of increased
lengths, find the point that fills it in the successive roundings until
the slot for $x$ is found.

All perfectly rounded $\mu$ have a curious property: both $m(x)=
\mu(x)/\mu(\{x\})$ and $\|m(x)\|=-\log\mu(\{x\})$ are always integers,
making $\mu(x)={.}m(x)\in Q_2$, a fraction expressed by bits of $m(x)$
following the dot. So $m$ is quite uniformly distributed:
$2k\mu(m^{-1}(k))\in[1,2]$ for $k\in m(\mathbb{N})$. It~is also
computable in polynomial time, as $m^{-1}$ is (by binary search). So, we
can use $m(x)$ as an alternative representation for $x$, in which the
distribution $\mu$ is remarkably uniform.

Simple distributions are not normally general enough. They may be the
ultimate source of the information in the instances $x$ of our problems,
but the original information $r$ is transformed into~$x$ by some process
$A$ that may itself be something like a one-way function. We can assume
that~$A$ is an algorithm with a reasonable time bound, but not that its
output distribution is simple. Such distributions are called
\emph{samplable}. In \cite{samp}, samplable distributions are dealt with
in a similar manner as with those in this section, though through a
different trick.

\section{Completeness}

\subsection{Complete Distributions and Inverters}

Given a function to invert, how would one generate hard instances? There
are two aspects of the problem. The first is achieving a significant
probability of generation of hard instances. The second is keeping the
probability of easy instances negligible. A number of reductions (with
various limitations) exist between these tasks. Let us restrict our
attention to the first one.

First, note that Lemma~\ref{rnd} enumerates all distributions computable
in time $t(x)$, preserving $t$ within a linear factor. Thus, we can
generate the largest of all these distributions by adding them up with
summable coefficients, say, $1/i^2$. This distribution will be complete
for TIME$(t(x))$ and belong to TIME$(t(x)\|x\|)$. A nice alternative
would be to combine all complexities by translating high times into
small probabilities, similarly to Section~\ref {las}. Instead, we will
switch to samplable distributions directly.

\begin{definition}
Distributions generated by algorithms in L without input are called
\emph{samplable}. If the algorithm has inputs, they are treated as a
parameter for a \emph{family of samplable distributions}.
\end{definition}

Usual versions of this definitions are broader, allowing a class of
algorithms closer to LV(P) and generation of polynomially larger
distributions; we limit the generators to L for greater precision.

\begin{proposition}\label{opt}
There exists a complete, i.e., largest up to constant factors, family of
samplable distributions. \end{proposition}

The lemma follows if we note that L is enumerable and, thus, the
complete distribution can be obtained by choosing members of L at random
and running them. The generator of this distribution spends $O(1)$
average time per run and has the greatest in L (up to a constant factor)
capacity for generating surprises. (Compared to LV(P) algorithms, its
chance of a nasty surprise may be polynomially smaller.)

Since the complete distribution makes sense only to within a constant
factor, it is robust in logarithmic scale only and defines an objective
hardness of ``hitting'' a set $X$ given $x$.

\begin{notation}
By $\mathop{\rm Kl}(X/x)$, we denote the $-\log$ of probability of a set
$X$ under the complete distribution family parameterized by $x$.
\end{notation}

As often happens, a strong attack tool helps inventing a strong defense.
Optimal searches were noted, e.g., in \cite{search,laws,bnt} but here we
get a nice version for free. The complete generator of hard instances
can be used as an optimal algorithm for solving them. Since algorithms
in L combine time into probability, their performance is measured by
their chance of success. The generator of an optimal distribution family
(parameterized by the instance $x$) has the highest, up to a constant
factor, chance $1/S(f/x)=2^{-\mathop{\rm Kl}(f^{-1}(x)/x)}$ of
generating solutions. Its minus logarithm $s(f/x)$ measures the hardness
of each individual instance $x$ and is called its \emph{security}. The
generator takes an $O(1)$ average time per run and succeeds in expected
$S(f/x)$ runs. No other method can~do~better.

\paragraph{Open problem.}
The constant factor in the optimal inversion algorithm may be
arbitrarily large. It is unknown whether it can be limited to a fixed
constant (say, 10) independent of the competing algorithm for
sufficiently large instances.

\subsection{Inversion Problems and OWFs}

A complete distribution achieves about as high probability of hard
instances as possible. Using it makes the choice of a hard-to-invert
function easy: all NP-complete functions would do similarly well.
However, the interesting goal is usually to find a function that is
tough for some standard, say, uniform distribution of instances.
Serendipitously, Lemma~\ref{rnd} transforms any P-time distribution into
a uniform one via an appropriate encoding. Combined with this encoding,
any NP-complete function joins those hardest to invert. However, the
encoding destroys the function's appeal, so the problem of combining a
nice function with a nice distribution remains.

Of course, while many functions seem hard, none are proven to be such.
In \cite{rp,samp,color,G90,VR,JW}, a~number of combinatorial and
algebraic problems are proved to be complete on average with uniform
distributions, i.e., as hard as any inversion problem with samplable
distribution could be.

These results, however, do not quite yield owfs. The difference between
owf and complete on average inversion problems can be described in many
ways. The simplest one is to define owfs as hard on average problems of
inverting \emph{length-preserving} functions. In this case the choice
between picking at random a witness (crucial for owfs), or an instance,
becomes unimportant.

Indeed, each witness gives only one instance. So if length is preserved,
the uniform distribution restricted to solvable instances does not
exceed the one generated by uniformly distributed witnesses. On the
other hand, if the witnesses are mapped into much fewer hard instances,
they must have many siblings. Then, the function can be modified as
follows. Guess the logarithm $k$ of the number of siblings of the
witness $w$ and pick a random member $a$ of a universal hash family
$h_a(w)$. Output $f(w),k,a,h'_a(w)$, where $h'$ is $h$ truncated to $k$
bits. (These outputs, i.e. instances, are $k$ bits too long but can be
hashed into strings of the same length as the witnesses.) The extra
information in the output (if $k$ is guessed correctly) is nearly
random, and so makes inversion no easier. However, the siblings are
separated into small groups and the numbers (and, thus, uniform
probabilities) of hard instances and their witnesses become comparable.
The converse is also true:

\begin{proposition} Any owf with multimedian $V(k)$ of its security
$S(x)$ (for uniformly random $w$ and $x=f(w)$) can be transformed into a
length-preserving owf with a $1/O(k)$ fraction of instances that have
security polynomially related to $V$. \end{proposition}

First, the fraction of hard instances is boosted as described at the end
of Section~\ref{mult}. If the number of hard instances is much smaller
than that of their witnesses, the function can still be made
length-preserving without altering its hardness by separating the
siblings as in the previous paragraph. See \cite{samp} for more detailed
computations of the results of hashing owfs.

\subsection{Complete OWF: Tiling Expansion}

We now modify the Tiling Problem to create a complete combinatorial owf.
No such owf has been described yet, though \cite{ow} shows the existence
of an artificial complete owf. It would be nice to have several complete
owfs that are less artificial, i.e., intuitive to someone who does not
care to know the definition of computability. Below, such an example is
given as a seed. Hopefully, a~critical mass of such examples will be
achieved some day providing an arsenal for reductions to more popular
owf candidates to show their completeness.

\paragraph{NP versus owf completeness.}
It is a mystery why the industry of proving worst-case NP-completeness
of nice combinatorial problems is so successful. Of course, such
questions refer to art rather than science and so need not have a
definite answer. It seems important that several ``clean'' complete
combinatorial NP problems are readily available without awkward
complexities that plague \emph{deterministic} computation models. Yet,
\emph{average-case} completeness is only slowly overcoming this barrier.
And for complete \emph{owfs}, my appeal to produce even one clean
example remained unanswered for two decades.

A complete owf is easy to construct as a modification of a universal
Turing Machine (UTM). UTM computation, in turn, is easy to transform
into combinatorial objects (tiled squares, etc.). These objects are
stripped from the many technicalities needed to define a model of
computation. Indeed, these technicalities are involved in making
computation deterministic, which is unnecessary since the final relation
is intended to be nondeterministic anyway. The ``stripped'' versions
have simple combinatorial structure and appeal, which makes it possible
to find so many reductions to other mathematical objects.

This approach works for constructing average-complete NP problems. It
fails, however, to preserve length which is needed for constructing
complete owfs. (In some other definitions, length-preservation is
replaced with other requirements, but these, too, are destroyed by the
above construction.) Here, I use simple tools, such as the concept of
expansion, to try to make a first step to overcome these problems. The
construction preserves input length and retains clean combinatorial
structure of tiling. I hope that this first step could be used as a
master-problem for reductions to other nice owfs.

 \def\tilet {Tiles: unit squares with a letter at each corner;\\
             may be joined if the letters match.}
 \def\expan {Expansion: maximal tile-by-tile {\bf unique} (using given
tiles)\\ extension of a partial tiling of a square with marked border.}

\newlength\tilel\setlength\tilel{\columnwidth}
\newlength\tbx\setlength\tbx{10ex}
\addtolength\tilel{-\tbx}\addtolength\tilel{-1em}
\medskip\noindent\fbox{\parbox{\tilel}{\tilet\\ \expan}
\parbox{\tbx}{\mbox
 {\fbox{\parbox{3ex} {\makebox[3ex]{a\hfill x} \makebox[3ex]{e\hfill
	 r}}}
  \fbox{\parbox{3ex} {\makebox[3ex]{x\hfill c} \makebox[3ex]{r\hfill
	z}}}}
  \par\vspace {.5ex} \mbox
 {\fbox{\parbox{3ex} {\makebox[3ex]{e\hfill r} \makebox[3ex]{n\hfill
	 s}}}
  \fbox{\parbox{3ex} {\makebox[3ex]{r\hfill z} \makebox[3ex]{s\hfill
	z}}}}}}

\begin{definition}
\emph{Tiling Expansion} is the following function: Expand a given top
line of tiles to a square using a given set of \emph{permitted tiles};
output the bottom line and the permitted tiles. \end{definition}

\begin{theorem}
Tiling Expansion is a owf if and only if owfs exist. \end{theorem}

\paragraph{The reduction.}
We start with a UTM, add a time counter that aborts after, say, $n^2$
steps. We~preserve a copy of the program prefix and force it, as well as
the input length, on the output. This produces a complete
``computational'' length-preserving owf. Now, we reduce the computation
of the UTM, modified as above, to Tiling in a standard way. We add a
special \emph{border symbol} and restrict the tiles so that it can
combine only with the input or output alphabets (of equal size), or~with
the \emph{end-tape} symbol, or the state initiating the computation,
depending on the sides of the tile. The expansion concept does the rest.

Tiling is a nice combinatorial entity, but it lacks determinism. This
demands specifying all tiles in the square, which ruins the length
preservation. Requiring the set of tiles to force deterministic
computation would involve awkward definitions unlikely to inspire
connections with noncomputational problems, reductions to which is the
ultimate goal. Instead, expansion allows an arbitrary set of tiles but
permits the extension of a partially tiled square only at places where
such possible extension is unique. In this way, some partially tiled
squares can be completed one tile at a time; others get stuck. This
process is inefficient, but efficiency loss (small in parallel models)
is not crucial here.

It remains an interesting problem to reduce this owf to other nice
combinatorial or algebraic owfs thus proving their completeness.

The raw owfs, however, are hard to use. While many results, such as,
e.g., pseudo-random generators, require no other assumptions, their
constructions destroy efficiency almost entirely. To~be useful, owfs
need other properties, e.g., low Renyi entropy. This low entropy is also
required for transforming weak owfs into strong ones (while preserving
security, as in \cite{expnd}), and for other tasks. The following note
suggests a way that may achieve this. (Its $f(x)+ax$ can be replaced
with other hashings.)

\begin{remark}
Inputs of $g(a,x)=(a,f(x)+ax)$ have $\le1$ siblings on average for any
length-preserving~$f$ and $a,x\in$GF$_{2^{\|x\|}}$. \end{remark}

\begin{conjecture}
The above $g$ is one-way, for any owf $f$, and has the same (within a
polynomial factor) security. \end{conjecture}

\end{document}